you invested $21 020 part of it in a stock that paid 12% annual interest. The rest of the money suffered a 5% loss. It the total annual income from both investments was $1754 how much was invested at each rate? solve math problem
Answer by richwmiller(17219)

Total amount of money invested: $21020
x+y=21020,
Total yearly interest for the two accounts is: $1754
0.12*x+-0.05*y=1754
x=21020-y
Substitute for x
0.12*(21020-y)+-0.05*y=1754
Multiply out
2522.4-0.12*y+-0.05*y=1754
Combine like terms.
-0.17*y=-768.4
Isolate y
y=$4520.00 at -5% lost $226
x=21020-y
Calculate x
x=$16500.00 at 12% earned $1980 interest
Check
0.12*16500-0.05*4520=1754
1980-226=1754
1754=1754
codeint1